# Generating variable $\hbar$ and $c$ via Fujii–Wetterich model in curved spacetimes

### v1

@@ A Higgs–scalon coupling model—Consider the full action
which is proportional to $\chi_{*}$. In a local reference frame tangent to the manifold at the point $x^{*}$, the action for the matter sector (excluding the Higgs excitation above the vacuum) becomes
@@
$\left(i\gamma^{\mu}\partial_{\mu}+\sqrt{\alpha}\,\gamma^{\mu}A_{\mu}+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\right)\psi$ $=0$ (15)
@@
$\left(i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar c}}\,\gamma^{\mu}A_{\mu}+m\,\frac{c}{\hbar}\right)\psi=0\,.$ (17)
@@
In Eq. (17), the parameters $e$ and $m$ represent the $U(1)$ gauge charge and the mass of the electron, respectively. As they are intrinsic properties of the electron, we then require them to be parameters rather than fields  [^7] We must emphasize that $e$ and $m$ can ‘run’ in the RG flow when radiative corrections related to the matter action described by Eq. (14) are included.. That is to say, we condition that they are independent of $\chi_{*}$. By comparing Eqs. (15)–(16) against Eqs. (17)–(18), we identify
@@
These identifications then require that
@@
Here, the subscript * is used to indicate the dependence of $\hbar$ and $c$ as functions of $\chi$. It should be noted that the (Newton) gravitational constant in independent of $\chi_{*}$ because $G:=\frac{c_{*}^{3}}{\hbar_{*}\chi_{*}^{2}}=1$.
Conversely, in the tangent frame covering an open set around the point $x^{*}$, the action in Eq. (14) can be recast as an action of Quantum Electrodynamics (QED) with a Lagrangian expressed in terms of $\hbar_{*}$, $c_{*}$, $m$, and $e$:
@@
Therefore, each open set vicinity of a given point $x^{*}$ on the manifold is equipped with a replica of QED action (22), operating with an effective quantum of action $h_{*}$ and an effective speed of light $c_{*}$, both of which are determined by the prevailing value $\chi_{*}$ of the background scalon field. As a component of the gravitational sector, $\chi_{*}$ can vary across the manifold, leading to corresponding variations in $\hbar_{*}$ and $c_{*}$ throughout the manifold.
In each such open set, the effective speed of light $c_{*}$ governs the propagation of the (massless  [^8] Note that only the discrete $\mathbb{Z}_{2}$ symmetry (10) of the Higgs singlet is (spontaneously) broken. The $U(1)$ gauge symmetry in Eqs. (8) and (9) remains unbroken, leaving the gauge vector boson $A_{\mu}$ massless.) gauge vector field $A_{\mu}$, whereas the effective quantum of action $\hbar_{*}$ regulates the quantization of the fields $\psi$, $A_{\mu}$, and $\Phi$. We should note that our mechanism to generate variable $\hbar$ and $c$ is readily applicable to the generalization of $\mathcal{L}_{\text{mat}}$ in Eq. (7) to the full GWS model of particle physics, where the gauge group is enlarged to $SU(3)\times SU(2)\times U(1)$. This extension is relatively straightforward and will be presented elsewhere [13].
@@
Here, the matrices are $\alpha^{k}:=\gamma^{0}\gamma^{k}$ and $\beta:=\gamma^{0}$. The time evolution of $\psi(t)$ within the open set vicinity of $x^{*}$ thus depends on the behavior of $\hbar_{*}$ with regard to $\chi_{*}$. This crucial connection between $\hbar_{*}$ and the evolution rate of quantum states leads to a concrete prediction, which we shall elaborate in subsequent sections.
A prediction: Anisotropic scaling in the clock rate—The QED action given in Eq. (14) possesses a dilatation symmetry, meaning it remains invariant under the transformation
@@
The inverse of the prevailing value $\chi_{*}^{-1}$ of the scalon field $\chi$ at the point $x^{*}$ thus sets the length scale for physical processes within the open set surrounding $x^{*}$, i.e.
@@
This is why we have referred to the field $\chi$ as a ‘scalon’.
However, due to the dependence of the speed of light on $\chi_{*}$ as derived in Eq. (21), the time scale for physical processes in the open set surrounding $x^{*}$ exhibits an anomalous behavior
@@
To verify the behavior of the length and time scales with respect to $\chi_{*}$, we shall consider the Hydrogen atom as an example. The Bohr radius is
@@
which is consistent with Eq. (26). The (relativistic) energy level of electron in a quantum state $\left|n,\,j\right\rangle$ of a hydrogen atom is a well-established result [36]  [^9] In the taylor expansion $E_{n}^{j}=mc_{*}^{2}-\frac{\alpha^{2}}{2n^{2}}mc_{*}^{2}+\mathcal{O}\left(\alpha^{4}\right)$, the leading term is the rest-mass energy and the sub-leading term recovers the energy level known in the non-relativistic limit for the Hydrogen atom.
@@
in which $\mathcal{N}_{n}^{j}\!:=\!{\scriptstyle{\scriptstyle\left(1+\alpha^{2}\left(n-j-\frac{1}{2}+\sqrt{\left(j+\frac{1}{2}\right)^{2}-\alpha^{2}}\right)^{-2}\right)^{-1/2}}}$ with $n=1,2,3,...$ and $j=\frac{1}{2},\frac{3}{2},\frac{5}{2},...$. Note that energy is proportional to $\chi_{*}$. The groundstate $\left|n=1,\,j=1/2\right\rangle$ and the excited state $\left|n=2,\,j=3/2\right\rangle$ have energy
@@
$\nu$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{\hbar_{*}}$ (33)
$=\frac{f\mu}{\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}^{3/2}\,.$ (34)
@@
which is in perfect alignment with Eq. (27). This time scaling result is generic since the time scale of the electron in a quantum state $\left|n,\,j\right\rangle$ can be deduced via Ehrenfest’s theorem, $\left\langle v_{n}^{j}\right\rangle=\frac{1}{m}\left\langle p_{n}^{j}\right\rangle=\frac{1}{mc_{*}}\left(\left(E_{n}^{j}\right)^{2}-m^{2}c_{*}^{4}\right)^{1/2}$. For example, considering the groundstate of the hydrogen atom, the typical time scale for the electron to complete its ‘orbit’ is
$\tau=\frac{2\pi\,a_{B}}{\left\langle v_{n=1}^{j=1/2}\right\rangle}=\frac{2\pi\sqrt{\lambda}}{\alpha^{2}f\mu}\,\chi_{*}^{-3/2}$ (36)
which is in full conformity with Eqs. (27) and (35).
The time scaling (27) implies that the evolution rate of a clock—regardless of whether it is a mechanical clock, an electronic clock, or an atomic clock—varies in spacetime, as a function of $\chi$, in an anisotropic fashion. In principle, this effect can be measured experimentally: Prepare two identical clocks at a location $A$. Keep one clock at location $A$ and send the second clock to a location $B$. Suppose that the background scalon field has different values $\chi_{A}$ and $\chi_{B}$ at the two locations $A$ and $B$, respectively. At their respective locations, the clocks would run at different rates per
@@
This predicted time dilation effect, while similar to the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric, arises from the dependence of the clock rate on the scalon field $\chi$, viz. Eq. (37). This new phenomenon is distinct and shall be referred to as a “time dilation effect of the Third kind”, to be investigated elsewhere [16].  [^10] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer our predicted time dilation effect as “the Third kind”.
Decay rate due to spontaneous emission—The prediction that we have just made can be validated with a different setup. Reconsider the Hydrogen atom. Induced by electric dipole perturbation in the vacuum, the electron in the excited state $\left|n=2,\,j=3/2\right\rangle$ can spontaneously transition to the groundstate $\left|n=1,\,j=1/2\right\rangle$ (allowable by the selection rule $\Delta j=\pm 1$). According to quantum mechanics [37], the decay rate (i.e. Einstein’s coefficient) is known to be
@@
where we have made the subscripts * in $\hbar$ and $c$ explicit. The angular frequency (with $h_{*}:=2\pi\hbar_{*}$) is
@@
Thus, the lifetime of this unstable quantum system behaves as
@@
which perfectly conforms with Eq. (27). Consequently, the clock involved in our experiment proposed in the preceding section can be any generic timekeeping device. It can be a mechanical clock, an electronic clock, an atomic clock  [^11] The Hafele–Keating experiment [38, 39] was carried out using atomic clocks., or even a radioactive quantum system.
Comparing with the Fujii–Wetterich scheme—We must stress that the dependence of the clock rate on $\chi_{*}$ has been documented in the works of Fujii and Wetterich [7, 8, 9, 10, 11, 12], although it was not a focal point in their analysis; see Footnote 2. However, their findings diverge from ours, particularly in how the clock rate scales with $\chi_{*}$. Specifically, they predicted a relation
@@
Consequently, the gravitational constant varies as $G_{\text{FW}}:=\frac{c_{\text{FW}}^{3}}{\hbar_{\text{FW}}\chi_{*}^{2}}=\chi_{*}^{-2}$. Instead of Eq. (23), the time evolution of the electron wavefunction in their scheme is governed by
@@
with the Hamiltonian $\hat{H}(t)$ given in Eq. (24). Here, $\hbar_{FW}$ is constant (equal to $1$). Therefore, the time scale for evolution in their scheme scales as
@@
This result can be verified with the concrete examples we used previously. In the Fujii–Wetterich scheme, the frequency of emitted photon from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ is
$\nu_{\text{FW}}$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{\hbar_{\text{FW}}}$ (47)
$=\frac{f\mu}{\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}$ (48)
@@
Therefore, despite having an identical matter action, Eq. (14), our scheme and the Fujii–Wetterich scheme result in decisively different predictions for the behavior of the clock rate. Future technologies may be able to distinguish the two predictions, and hence the validity of our scheme versus that of the Fujii–Wetterich scheme.
Advantages of our mechanism—Our approach offers two distinct advantages over the Fujii–Wetterich treatment [7, 8, 9, 10, 11, 12]:
  1. 1.
    The Fujii–Wetterich scheme treats particle masses and charges at disparity: while masses are promoted to fields dependent on $\chi_{*}$, charges remain as fixed parameters (see Eq. (44)). Given that both masses and charges are intrinsic properties of particles, our scheme treats them on equal footing: particle masses and charges are parameters but not fields. (Note: in our scheme, $m$ and $e$ can ‘run’ in the RG flow, but they are independent of the background scalon $\chi$.)
  2. 2.
    The Fujii–Wetterich scheme only applies to massive particles, leaving massless particles unaffected. In contrast, the varying $c$ and $\hbar$ in our scheme impact all particles—massive and massless alike. Specifically, the variation in $c$ influences the propagation of (massless) photon in an expanding cosmic void, a feature absent in the Fujii–Wetterich scheme.
The anomalous $3/2$–exponent in our time scaling (27) is a novel discovery, leading to new physics with a unique set of previously unexplored phenomena and predictions. Specifically, it induces the variability of $c$ in spacetime, per $c_{*}:=l/\tau\propto\chi_{*}^{1/2}$, causing lightwaves traveling through the expanding cosmic void to undergo an additional refraction effect. This effect thence necessitates a reanalysis of the Pantheon Catalog of Type Ia supernovae, a task we will address in Section II.
II Phenomenology of VSL
This section focuses on our application of variable speed of light (VSL) for late-time cosmology, bypassing the need for dark energy (DE). The detailed work, including full technical elaboration, is presented in two companion papers, Refs. [14, 15].
I. VSL cosmography and cosmology—Our starting point is the generalized Friedmann–Lemaître–Robertson–Walker (FLRW) metric for the expanding cosmic void (with $\kappa=\{1,0,-1\}$)
$ds^{2}$ $=c^{2}\,dt^{2}-a^{2}(t)\,\left(\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\right)$ (50)

### v2

@@ III A new mechanism to generate variable $\hbar$ and $c$
which is proportional to $\chi_{*}$. In a local reference frame tangent to the manifold at the point $x^{*}$ (i.e, an inertial frame at $x^{*}$), the action for the matter sector (excluding the Higgs excitation above the vacuum) becomes  [^6] If we adopt the analysis presented in Ref. [133], the Higgs VEV $v(x):=\left\langle\Phi(x)\right\rangle$ would obey the equation $-2\square v+\mu^{2}\chi^{2}v-\lambda v^{3}=0$. However, we treat $v$ as a slowly varying background field; under this condition, the projection of the matter sector onto the local inertial frame, viz. Eq. (14), is justified.
@@
$\biggl{(}i\gamma^{\mu}\partial_{\mu}+\sqrt{\alpha}\,\gamma^{\mu}A_{\mu}+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\biggr{)}\,\psi$ $=0$ (15)
@@
$\biggl{(}i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar c}}\,\gamma^{\mu}A_{\mu}+m\,\frac{c}{\hbar}\biggr{)}\,\psi=0\,.$ (17)
@@
In Eq. (17), the parameters $e$ and $m$ represent the $U(1)$ gauge charge and inertial mass of the electron, respectively. As they are intrinsic properties of the electron, we then require them to be parameters rather than fields, namely, they are independent of $\chi_{*}$. (Note: We must emphasize that $e$ and $m$ are not constants, but they can ‘run’ in the renormalization group flow when radiative corrections involving $\psi$ and $A_{\mu}$ are included.) By comparing Eqs. (15)–(16) against Eqs. (17)–(18), we can identify
@@
These identities then lead to
@@
Here, the subscript * in $\hbar_{*}$ and $c_{*}$ signifies the dependence of $\hbar$ and $c$ as functions of $\chi$. It should be noted that Newton’s constant in independent of $\chi_{*}$ because $G:=\frac{c_{*}^{3}}{\hbar_{*}\chi_{*}^{2}}=1$.
Conversely, in the tangent—i.e., inertial—frame at the point $x^{*}$, the action in Eq. (14) can be recast as an action of Quantum Electrodynamics (QED) with a Lagrangian expressed in terms of $\hbar_{*}$, $c_{*}$, $m$, and $e$, viz.
@@
Therefore, each open set enclosing a given point $x^{*}$ on the manifold is equipped with a replica of the QED action (22), operating with an effective Planck constant $h_{*}$ and an effective speed of light $c_{*}$. Both of these effective parameters are determined by the prevailing value $\chi_{*}$ of the background dilaton field, per Eq. (21). As a component of the gravitational sector, $\chi_{*}$ can vary across the manifold, leading to corresponding variations in $\hbar_{*}$ and $c_{*}$ throughout the manifold.
Within each open set, the effective speed of light $c_{*}$ governs the propagation of the (massless  [^7] Note: Only the discrete $\mathbb{Z}_{2}$ symmetry (11) of the Higgs singlet is broken. The local $U(1)$ gauge symmetry in Eqs. (9) and (10) remains unbroken, leaving the gauge vector boson $A_{\mu}$ massless.) gauge vector field $A_{\mu}$, whereas the effective Planck constant $\hbar_{*}$ regulates the quantization of the fields $\psi$, $A_{\mu}$, and $\Phi$. We should note that our mechanism for generating variable $\hbar$ and $c$ is readily applicable to the generalization of $\mathcal{L}_{\text{mat}}$ in Eq. (8) to the full GWS model of particle physics, where the gauge group is enlarged to $SU(3)\times SU(2)\times U(1)$. This extension is a relatively straightforward exercise. A simplified version of our mechanism is also provided in Ref. [134] where we allowed the dilaton to interact directly with the fermion field in the form $\chi\,\bar{\psi}\psi$. We obtained results that are essentially identical to those presented above.
@@
Here, the matrices are $\alpha^{k}:=\gamma^{0}\gamma^{k}$ and $\beta:=\gamma^{0}$. The time evolution of $\psi(t)$ within the open set enclosing $x^{*}$ thus depends on the behavior of $\hbar_{*}$ with regard to $\chi_{*}$. This crucial connection between $\hbar_{*}$ and the evolution rate of quantum states leads to a concrete prediction, which we shall elaborate in the subsequent sections.
IV A prediction: Anisotropic scaling in the clock rate
The effective QED action given in Eq. (14) possesses a dilatation symmetry; namely, it remains invariant under the transformation
@@ IV A prediction: Anisotropic scaling in the clock rate
The prevailing value of the dilaton field $\chi$ at the point $x^{*}$ thus sets the length scale for a given physical process within the open set surrounding $x^{*}$, i.e.
@@
justifying the term “dilaton” for $\chi$. However, due to the dependence of $c$ on $\chi$ as given in Eq. (21), the timescale for the physical process in the open set exhibits an anisotropic behavior
@@
To illustrate the behavior of the length and time scales with respect to $\chi_{*}$, we will consider the Hydrogen atom as an example. The Bohr radius is
@@
which is consistent with Eq. (26). The energy level of (relativistic) electron in a quantum state $\left|n,\,j\right\rangle$ of a hydrogen atom is a well-established result [135]
@@
in which $\mathcal{N}_{n}^{j}\!:=\!{\scriptstyle{\scriptstyle\left(1+\alpha^{2}\left(n-j-\frac{1}{2}+\sqrt{\left(j+\frac{1}{2}\right)^{2}-\alpha^{2}}\right)^{-2}\right)^{-1/2}}}$ with $n=1,2,3,...$ and $j=\frac{1}{2},\frac{3}{2},\frac{5}{2},...$. Note that energy is proportional to $\chi_{*}$. The groundstate $\left|n=1,\,j=1/2\right\rangle$ and the excited state $\left|n=2,\,j=3/2\right\rangle$ have energy levels
@@
$\nu$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{2\pi\,\hbar_{*}}$ (33)
$=\frac{f\mu}{2\pi\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}^{3/2}\,.$ (34)
@@
which is in perfect agreement with Eq. (27).
The time scaling (27) implies that the evolution rate of a clock—regardless of whether it is a mechanical clock, an electronic clock, or an atomic clock—varies in spacetime, as a function of $\chi$, in an anisotropic fashion. In principle, this effect can be measured experimentally: Prepare two identical clocks at a location $A$. Keep one clock at location $A$ and send the second clock to a location $B$. Suppose that the background dilaton field has different values $\chi_{A}$ and $\chi_{B}$ at the two locations $A$ and $B$, respectively. At their respective locations, the clocks would run at different rates per
@@
This predicted time dilation effect differs from the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric, arises from the dependence of the clock rate on the dilaton field $\chi$, viz. Eq. (36). This new phenomenon is distinct and shall be referred to as a “time dilation effect of the Third kind”, to be investigated in future work.  [^8] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer our predicted time dilation effect as “the Third kind”.
Decay rate of unstable quantum systems
The prediction that we have just made can be validated via a different setup. Reconsider the Hydrogen atom. Induced by electric dipole perturbation in the vacuum, the electron in the excited state $\left|n=2,\,j=3/2\right\rangle$ can spontaneously transition to the groundstate $\left|n=1,\,j=1/2\right\rangle$ (allowed by the selection rule $\Delta j=\pm 1$). According to quantum mechanics [136], the decay rate (i.e. Einstein’s coefficient) is known to be
@@
where the effective $\hbar_{*}$ and $c_{*}$ are made explicit. The angular frequency (with $h_{*}:=2\pi\hbar_{*}$) is
@@
Thus, the lifetime of the decay process for this unstable quantum system behaves as
@@
which perfectly conforms with Eq. (27). Consequently, the clock involved in our experiment proposed earlier in this section can be any generic timekeeping device. It can be a mechanical clock, an electronic clock, an atomic clock, or even a radioactive quantum system.  [^9] The Hafele–Keating experiment [137, 138] was carried out using atomic clocks.
V Comparison of our VSL scheme with the Fujii–Wetterich scheme
We must stress that the dependence of the clock rate on $\chi_{*}$ has been documented in the works of Fujii and Wetterich [18, 12, 13, 14, 15, 16], although it was not a focal point in their analysis; see Footnote 2. However, their findings diverge from ours, particularly in how the clock rate scales with $\chi_{*}$. Specifically, they predicted a relation
@@ V Comparison of our VSL scheme with the Fujii–Wetterich scheme
Consequently, Newton’s “constant” varies as $G_{\text{FW}}:=\frac{c_{\text{FW}}^{3}}{\hbar_{\text{FW}}\chi_{*}^{2}}=\chi_{*}^{-2}$. Instead of Eq. (23), the time evolution of the electron wavefunction in the FW scheme is governed by
@@
with the Hamiltonian $\hat{H}(t)$ given in Eq. (24). Here, $\hbar_{FW}$ is constant (equal to $1$). Therefore, the time scale for evolution in the FW scheme scales as
@@
This result can be verified with the concrete examples we used previously. In the FW scheme, the frequency of emitted photon from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ is
$\nu_{\text{FW}}$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{2\pi\,\hbar_{\text{FW}}}$ (46)
$=\frac{f\mu}{2\pi\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}$ (47)
@@
Therefore, despite having an identical matter action, Eq. (14), our scheme and the FW scheme are not physically equivalent. They result in decisively different predictions for the behavior of the clock rate. Future technologies may be able to distinguish the two predictions, and hence the validity of each scheme.
Properties of our mechanism
Our approach offers two distinct benefits:
1. The FW scheme treats inertial mass and gauge charge at disparity: while mass is promoted to fields dependent on $\chi_{*}$, charge remains as a fixed parameter (see Eq. (43)). Given that inertial mass and gauge charge are intrinsic properties of particles, our scheme treats them on equal footing: particle mass and charge are parameters but not fields. (Note: in our scheme, $m$ and $e$ can ‘run’ in the RG flow, but they are independent of the background dilaton $\chi$.)
2. The FW scheme only applies to massive particles, leaving massless particles unaffected. In contrast, the varying $c$ and $\hbar$ in our scheme impact all particles—massive and massless alike. Specifically, the variation in $c$ influences the propagation of (massless) photon in an expanding intergalactic space in cosmology, a feature absent in the FW scheme.
In conclusion of our derivation, the $3/2$–exponent in our time scaling (27) is a novel discovery, leading to new physics with a unique set of previously unexplored phenomena and predictions. Specifically, it induces the variability of the speed of light in spacetime, causing lightwaves traveling through the expanding universe to undergo an additional refraction effect. This effect thence necessitates a reanalysis of the Pantheon Catalog of Type Ia supernovae, a task we carry out in the next section.
VI Phenomenology of VSL
This section focuses on our application of variable speed of light (VSL) for late-time cosmology, bypassing the need for dark energy (DE). The detailed work containing full technicality is presented in Refs. [19].
VI.1 VSL cosmology and late-time cosmography
Our starting point is to generalize the Friedmann–Lemaître–Robertson–Walker (FLRW) metric for the cosmic scale factor (with $\kappa=\{1,0,-1\}$)
$ds^{2}$ $=c^{2}\,dt^{2}-a^{2}(t)\,\biggl{(}\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\biggr{)}$ (49)